Question 4: Cost structure estimation Fixed or variable? OF OV OF OV Cost account direct labor direct materials depreciation rent administrative staff salaries Production volume was 200 units Amount $4,000 $6,000 $6,000 $7,000 $7,000 OF OV a) Which cost structure estimation method should we use? O high-low method O account classification method b) Classify each cost account above as fixed or variable. Enter your answers in the table above c) Estimate the cost structure parameters: total FC $ unit VC= $ Write down the total cost equation: (e.g., if TC-500+2 volume, enter 500 in the first box and 2 in the second box) d) Predict the total cost of producing 250 units: volume

=========================

MANAGERIAL ACCOUNTING QUESTIONS

Please answer ALL in DETAILS to get upvote

Thanks!!!
Answer #1

(a)

Account classification method should be used here to classify costs into fixed costs and variable costs. High low method of cost classification is used when different output levels and total cost at different output levels is given.

(b)

Cost account Amount Fixed or variable
Direct labor $4,000 Variable
Direct material $6,000 Variable
Depreciation $6,000 Fixed
Rent $7,000 Fixed
Administrative staff salaries $7,000 Fixed

(c)

Total fixed cost = 6,000 + 7,000 + 7,000

= $20,000

Total variable cost = 4,000 + 6,000

= $10,000

Production volume = 200 units

Variable cost per unit = Total variable cost/Production volume

= 10,000/200

= $50

Total cost = 20,000 + 50 x Volume

(d)

Cost of producing 250 units = 20,000 + 50 x 250

= 20,000 + 12,500

= $32,500

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.